OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -16 when x = 4, find x when y = 32.

OpenStudy (dmndlife24):

Inverse variation is y = k/x When y = -16 and x = 4 we need to find the constant -16 = k/4 We find that k = -64 If we are asked to find x when y = 32 we can do that bc we have our constant 32 = -64/x 32x = -64 x = -2